OpenStudy (anonymous):

-4w^8 x (-6w^7) someone please help.

OpenStudy (anonymous):

i dont understand how to solve it

OpenStudy (anonymous):

\[-4\times (-6)w^{8+7}\] is a start

OpenStudy (anonymous):

(ax^n ) * (bx^m) = (a*b) x^(n+m)

OpenStudy (anonymous):

so then it would be, 24w^15?

OpenStudy (xxdragoonxx):



OpenStudy (anonymous):

oh so i was right lol

OpenStudy (anonymous):

somehow its wrong O.o

OpenStudy (anonymous):

idk

OpenStudy (xxdragoonxx):

It looks right to me.

OpenStudy (anonymous):

yeah, thats what im confused about like i know it was solved right... hmm